OpenStudy (anonymous):

a Last fall, a gardener planted 60 iris bulbs. She found that only 45 of the bulbs bloomed in the spring. Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms. How many of the bulbs should she plant next fall if she would like at least 52 to bloom?

OpenStudy (cwrw238):

45/60 = 3/4

OpenStudy (cwrw238):

second part = 52 * 4/3